# Randomness, Not Representation: The Unreliability of Evaluating Cultural Alignment in LLMs

### v1

@@ 5.2. Extrapolability
Next, to assess how often alignment with a culture on one issue predicts alignment on others, we used a three-phase methodology with the VSM dataset. First, we randomly divided the 75 countries into five equal groups, each assigned to one LLM. Second, we had each model complete the VSM survey for its assigned countries using a cultural steering prompt adapted from AlKhamissi et al. (2024) (see prompt in Table 8). Finally, we conducted progressive clustering analysis on both human and LLM-generated dimensional scores, incrementally including dimensions (n=1 to n=5) to test whether alignment on some dimensions predicts alignment on others. To evaluate clustering quality, we used the Adjusted Rand Index (ARI), which measures the similarity between two clusterings, scaled to range between 0 and 1 (scikit-learn developers, 2025), where higher values indicate stronger agreement between clusters (e.g., more extrapolability).
@@
Extrapolation based on a small number of cultural dimensions is unreliable. Figure 7 reveals that for small numbers of observed dimensions, extrapolating to others results in errors similar to the random chance baseline. However, as substantially more dimensions are considered, the ARI increases above $0.8$. This occurs for LLMs and humans alike. However, LLM predictions show lower disagreement than human country data, particularly at higher dimensions.
@@
The validity of extrapolation is highly sensitive to individual cultural dimensions. Figure 7 demonstrates that different cultural dimensions have varying impacts on clustering patterns. The ARI differences reveal how each dimension contributes to or detracts from cultural groupings when compared to established country clusters. Indulgence/Restraint shows the strongest positive contribution, indicating it reinforces expected cultural groupings, while Masculinity/Femininity demonstrates the strongest negative impact, suggesting it creates unexpected cultural associations that diverge from traditional groupings.
@@ 5.3. Steerability
To evaluate whether prompting can reliably steer LLMs to adopt specific cultural viewpoints, we compared LLM responses against human responses from 15 countries using the GQA dataset. We implemented and tested two distinct prompt steering approaches. First, we adapted the instruction-based prompt from AlKhamissi et al. (2024) with minimal modifications from Section 5.2. (See Appendix A for the prompt). We then used this prompt to evaluate LLM performance across all 180 questions. Second, we used DsPY (Khattab et al., 2023) with the MIPROv2 optimizer to generate and optimize a few-shot prompt using 90 questions and human answer pairs, conducting subsequent analysis on the remaining 90 questions.
LLMs exhibit erratic, un-humanlike responses to GQA questions under attempted prompt steering. Figure 8 shows t-SNE embeddings of human and model responses for both experiments. In both cases, human responses from different countries (blue) cluster closely together, while model responses show erratic, un-humanlike patterns, failing to cluster with human responses. Quantitatively, the ratios of average distances between LLM response pairs versus human response pairs were 6.14 and 1.94 for the AlKhamissi et al. (2024) and DSPy (Khattab et al., 2023) prompts respectively [^4] If we calculate the same ratio using each of the models separately, the minimum ratios are still well above 1.0 at 5.66 and 1.38 respectively.. Permutation tests against the null hypothesis that human-model response distances were identically distributed returned $p=0.0$ for both the AlKhamissi et al. (2024) and DsPY (Khattab et al., 2023) prompting methods.
Figure 8. LLMs fail to align with human views, let alone those of specific cultures under prompt steering. t-SNE embeddings of LLMs prompted to align with different human cultures using prompts from AlKhamissi et al. (2024) and optimized prompts from DSPy (Khattab et al., 2023). Human responses from different nations cluster together while LLMs exhibit erratic, un-humanlike responses. Permutation tests against the null hypothesis that human-model response distances were identically distributed returned $p=0.0$ for both the AlKhamissi et al. (2024) and DsPY (Khattab et al., 2023) prompting methods.
@@ 6. Manipulating LLM Evaluations with Forced Binary Choices: a Case Study
Throughout Section 5, we have shown that evaluations of LLMs’ cultural preferences can be surprisingly unstable, non-extrapolable, and unsteerable. An overarching theme of these results is that an LLM’s apparent cultural preferences in a narrow evaluation context can be misleading about its behaviors in other contexts. This prompts concerns about whether it is possible to strategically design experiments and cherry-pick results to paint an arbitrary picture of an LLM’s cultural preferences. Here, we present a case study in evaluation manipulation by showing that using Likert scales with versus without a ‘neutral’ option can suggest very different conclusions.
@@
Recent work by Mazeika et al. (2025) introduces the concept of “utility engineering” to analyze emergent value systems in large language models. Their experiments rely on forced binary choice prompts, requiring LLMs to select between two distinct options without allowing for expressions of neutrality. For example, they present LLMs with binary choices about whether they favor saving X human lives from country A or Y human lives from country B. The experimental results suggest that models like GPT-4o appear to value human lives differently based on nationality. However, we hypothesized that this particular finding may have been related to the lack of a neutral option in the binary choices presented to the LLMs. To test this, we conducted a comparative case study examining how preference elicitation format affects reported values.
6.1. Methodology
To evaluate the impact of forced binary choice on LLMs’ apparent preferences, we compared the preferences expressed with and without a neutral option. We implemented two rating formats:
  • Standard Likert Scale (odd number of options): A 5-point scale including an explicit neutral midpoint. This allows models to express that they value lives from different countries equally.
  • Forced Choice Likert Scale (even number of options): A 4-point scale excluding an explicit neutral midpoint. This forces models to express a preference that favors some lives over others.
See Appendix A for the complete prompts.
We tested these formats using the same set of countries examined in the exchange rate analysis from Mazeika et al. (2025). We generated all possible country pair permutations with bidirectional testing. We used GPT-4o with temperature 1.0 to ensure direct comparability with the original study’s findings. We repeated each comparison three times to control for potential stochasticity in model responses.
6.2. Results
@@
Figure 9. GPT-4o equally values human lives from different countries when it is able to choose neutrality between comparisons. However, when forced to express a preference (as in Mazeika et al. (2025)), it exhibits unequal values. Side-by-side comparison of country preferences on a 0-1 scale using the same country ordering. (Left) The 5-point Likert scale with a neutral option shows more uniform scores across countries. (Right) The 4-point forced-choice scale reveals more varied preferences, with a wider range of scores. Both visualizations use the same country order based on the forced-choice ranking to facilitate direct comparison.
We find significant differences in preference patterns across elicitation formats. In the 5-point Likert scale condition, GPT-4o selected the neutral midpoint (“No preference”) in 100% of country pair comparisons. This resulted in perfectly uniform normalized preference scores of 0.50 for all 11 countries (Figure 9).
The 4-point forced choice condition produced non-uniform country preferences (Figure 9). Without a neutral option available, we find, similarly to Mazeika et al. (2025), that the model displayed noticeable preferences for certain countries over others.
6.3. Methodological Implications
Our results demonstrate how evaluation design affects preference measurement in LLMs. The difference between GPT-4o’s uniform valuations with neutral options and its varied preferences under forced choice indicates that observed patterns can be specific to methodology rather than reflecting inherent model characteristics.
These findings provide context for interpreting Mazeika et al.’s (2025) exchange rate results, where they report that “GPT-4o places the value of Lives in the United States significantly below Lives in China, which it in turn ranks below Lives in Pakistan,” and suggest these represent “deeply ingrained biases” in the model. Our experiments with the same model reveal that these hierarchical preferences emerge primarily under forced-choice conditions that prevent the expression of equality. When allowed to express neutrality or evaluated through absolute importance ratings, GPT-4o consistently indicates equal valuation of human lives regardless of nationality, suggesting a more nuanced interpretation of the model’s apparent preferences.
@@ 6.3. Methodological Implications
This illustrates a critical limitation in extracting preferences from LLMs. Rather than revealing stable internal preferences, our findings show that LLM outputs are largely constructed responses to specific elicitation paradigms. Interpreting such outputs as evidence of inherent biases without examining methodological influence risks misattributing artifacts of evaluation design as properties of the model itself.
@@ 7. Discussion
Significance: It is appealing to assume that modern LLMs exhibit relatively stable, coherent, and steerable preferences. These three goals are central to the modern field of AI Alignment. Meanwhile, prior research on evaluating LLM cultural alignment has attempted to identify clear differences in how well they align with different cultures. However, we have found that state-of-the-art LLMs express surprisingly erratic cultural preferences. To the extent that LLMs are more ‘aligned’ to some cultures than others, it is in a nuanced and often context-specific way. As we have shown, small changes in evaluation methodology can lead to large changes in results. Our case study on human life valuation illustrates this clearly. GPT-4o uniformly values lives across countries when allowed to express neutrality, yet produces hierarchical preferences under forced-choice conditions. Ultimately, this cautions against drawing broad conclusions from narrow experiments, as methodological design choices may drive findings that appear to indicate inherent model biases. In particular, they suggest that simplistic evaluations, cherry-picking, and confirmation biases may contribute to an overly simplistic understanding of LLM cultural alignment.
@@
Limitations: The central theme of this work is about the risks of drawing broad conclusions from narrow evaluations. Similar to previous research, we also only conduct a limited set of experiments. It is clear that cultural stability, extrapolability, and steerability do not hold in general for state-of-the-art LLMs. Nonetheless, there might be narrow circumstances in which they do. For example, Benkler et al. (2023); Azzopardi and Moshfeghi (2024), and Jiang et al. (2024) all argue that they may hold in certain narrow cases. While cultural alignment evaluations suffer from fundamental limitations, our work should not be understood as showing that they are fundamentally invalid.
Future Work: Given the difficulties of cultural alignment evaluations in LLMs, future work could aim to improve models or evaluation strategies. Current models are fine-tuned to elicit approval from evaluators (e.g., with RLHF), but specifically fine-tuning them to be steerable and express consistent preferences could reduce the potential for existing evaluations to produce simplistic or misleading results. Meanwhile, improving the realism of evaluations offers another useful direction. LLMs (and humans alike) align with values in a way that is complex and context-dependent. Currently, we have a limited understanding of how cultural biases may affect LLM behaviors in subtle ways that are difficult to study with simple question/answer evaluations. Finally, it is reasonable to question the importance of evaluating preferences in LLMs entirely. Instead of asking what cultures they align with, it will also be useful to study how their deployments shift power and tangibly impact different communities in the real world (Kalluri, 2020).

### v2

@@ 5.2. Extrapolability
To test whether alignment with one culture on a narrow set of issues predicts alignment on others, we performed clustering analysis on cultural dimensions and measured consistency between partial and complete dimension sets. We used K-means to group countries based on their cultural values and evaluated similarity using the Adjusted Rand Index (ARI), which measures agreement between two clustering assignments on a scale from 0 to 1 (scikit-learn developers, 2025).
Our experimental procedure was as follows: First, we randomly divided 65 countries into five equal groups, each assigned to one of our five LLMs, to efficiently distribute computational resources while maintaining comprehensive coverage. Second, we had each model complete the VSM survey for its assigned countries using a cultural steering prompt adapted from AlKhamissi et al. (2024) (see Table 8).
For our analysis, we first established a reference clustering using all six Hofstede dimensions. We then systematically created new clusterings using every possible subset of dimensions (e.g., single dimensions, pairs, triplets, etc.). For each subset, we calculated the ARI between its resulting clustering and the reference clustering from all six dimensions. Higher ARI values indicate stronger agreement between clusterings (i.e., greater extrapolability).
To quantify the impact of individual dimensions, we compared the mean ARI of all subsets that included a specific dimension against those that excluded it. The resulting difference in mean ARI represents the dimension’s contribution to clustering consistency. Positive values indicate the dimension reinforces expected cultural groupings, while negative values suggest the dimension introduces different grouping patterns. This approach allowed us to test whether alignment on some dimensions predicts alignment on others and to identify which cultural aspects are most critical for accurate cultural extrapolation.
@@
Extrapolation based on a small number of cultural dimensions is unreliable. Figure 5 (Left) reveals that for small numbers of observed dimensions, extrapolating to others results in errors similar to the random chance baseline. However, as substantially more dimensions are considered, the ARI increases above $0.8$ for LLMs and humans alike. Nevertheless, LLM predictions show lower disagreement than human country data, particularly at higher dimensions.
@@
The validity of extrapolation is highly sensitive to individual cultural dimensions. Figure 5 (Right) demonstrates that different cultural dimensions have varying impacts on clustering. This shows that the validity of extrapolation can be sensitive to the specific cultural dimensions that are analyzed. For example, Indulgence/Restraint shows the strongest positive contribution, indicating it reinforces expected cultural groupings, while Masculinity/Femininity demonstrates the strongest negative impact, suggesting different patterns among these preferences compared to other dimensions.
Figure 5. Extrapolation across cultural dimensions is unreliable in humans and LLMs alike. The validity of extrapolation is highly sensitive to the geometry of individual cultural dimensions (Left) For humans and LLMs, extrapolability (as measured by clustering ARI) increases with the number of observed dimensions. However, it is near a random guess baseline for low numbers of observed dimensions. (Right) Different cultural dimensions have very different impacts on extrapolation between dimensions. For humans (blue), Indulgence/Restraint strengthens groupings while Masculinity/Femininity weakens them. For LLMs (orange), Uncertainty Avoidance Index strengthens while Long/Short Term orientation weakens the clustering. Hypothesis testing: */**∗ ∗/***∗ ∗ ∗ = $p<$0.05/0.01/0.001. (Left) For each group (Countries/LLM), the null hypothesis was that the clustering similarity between subsets of dimensions versus all dimensions could arise from random cluster assignments. (Right) For each group (Countries/LLM), the null hypothesis was that there is no difference in mean ARI scores between dimension subsets that include vs. exclude each dimension.
@@ 5.3. Steerability
To evaluate whether prompting can reliably steer LLMs to adopt specific cultural viewpoints, we compared LLM responses against human responses from 15 countries using the GQA dataset. We implemented and tested two distinct prompt steering approaches. First, we adapted the instruction-based prompt from AlKhamissi et al. (2024) with minimal modifications from Section 5.2. (See Appendix A for the prompt). We then used this prompt to evaluate LLM performance across all 180 questions. Second, we used DsPY (Khattab et al., 2023) with the MIPROv2 optimizer to generate and optimize a few-shot prompt using a 50% split of questions and human answer pairs for training and evaluation.
Our steerability analysis addresses whether prompt steering can make LLMs align with real human opinions from specific cultural contexts. Rather than simply measuring if prompting changes LLM behavior, we quantify how closely LLM responses resemble actual human responses from the target culture. To quantify this alignment, we computed Euclidean distances between normalized response vectors, where each element represents a Likert-scale response normalized to a [0,1] range. We measured the average pairwise distance between human responses and compared this to the average pairwise distance between LLM responses to determine whether responses cluster by country (indicating successful steerability) or by model architecture (indicating failed steerability). The T-SNE visualization illustrates these patterns, but our conclusions are based on the quantitative analysis of distances between survey responses.
LLMs exhibit erratic, un-humanlike responses to GQA questions under attempted prompt steering. Figure 6 shows t-SNE embeddings of human and model responses for both experiments. In both cases, human responses from different countries (blue) cluster closely together, while model responses show erratic, un-humanlike patterns, failing to cluster with human responses. Quantitatively, the ratios of average distances between LLM-human response pairs versus human-human response pairs were well above 1.0 at 6.14 and 1.94 for the AlKhamissi et al. (2024) and DSPy (Khattab et al., 2023) prompts respectively [^2] If we calculate the same ratio using each of the models separately, the minimum ratios are still well above 1.0 at 5.66 and 1.38, respectively.. This means that humans from other countries are better proxies for each other’s cultural preferences than LLMs, even when prompted to align with humans from any given country. Permutation tests against the null hypothesis that human-model response distances were identically distributed returned $p=0.0$ for both the AlKhamissi et al. (2024) and DsPY (Khattab et al., 2023) prompting methods.
Figure 6. LLMs fail to align with human views, let alone those of specific cultures under prompt steering. t-SNE embeddings of LLMs prompted to align with different human cultures using prompts from AlKhamissi et al. (2024) and optimized prompts from DSPy (Khattab et al., 2023). Human responses from different nations cluster together while LLMs exhibit erratic, un-humanlike responses. Permutation tests against the null hypothesis that human-human and human-model response distances were identically distributed returned $p=0.0$ for both the AlKhamissi et al. (2024) and DsPY (Khattab et al., 2023) prompting methods.
@@ 6. Manipulating LLM Evaluations with Forced Binary Choices: a Case Study
Throughout Section 5, we have shown that an LLM’s apparent cultural preferences in a narrow evaluation context can be misleading about its behaviors in other contexts. This raises concerns about whether it is possible to strategically design experiments or cherry-pick results to paint an arbitrary picture of an LLM’s cultural preferences. Here, we present a case study in evaluation manipulation by showing that using Likert scales with versus without a ‘neutral’ option can produce very different results.
@@
Recent work by Mazeika et al. (2025) proposed that LLMs develop emergent value systems that cause them to value different human lives differently. Their experiments rely on forced binary choice prompts, requiring LLMs to select between two distinct options without allowing for expressions of neutrality. For example, they present LLMs with binary choices about whether they favor saving X human lives from country A or Y human lives from country B. Their results suggest that models like GPT-4o appear to value human lives differently based on nationality. However, we hypothesized that this particular finding may have been related to the lack of a neutral option in the binary choices presented to the LLMs.
Methodology: To examine how forced binary-choice evaluations influence expressed preferences, we conducted an experiment comparing responses collected with and without a neutral response option. Specifically, we contrasted a standard 5-point Likert scale, including a neutral midpoint that allows models to indicate equal valuation of lives (see Table 9), against a forced-choice 4-point Likert scale excluding the neutral midpoint, forcing models to express a preference that favors some lives over others (see Table 10 for complete prompts).
We tested these formats using the same set of countries examined in the exchange rate analysis from Mazeika et al. (2025). For thoroughness, we generated all possible country-pair combinations and conducted comparisons bidirectionally (i.e., evaluating preferences in both directions for each country pair). To derive a single preference score for each country, we collected all pairwise comparisons where that country appeared and calculated the mean preference rating, appropriately inverting scores when necessary depending on the country’s position in each comparison. Ratings obtained from both scales were normalized to a 0–1 range and averaged across three repeated trials to account for stochasticity. All comparisons were performed using GPT-4o at temperature 1.0, matching the conditions from the original study.
@@
Figure 7. GPT-4o equally values human lives from different countries when it is able to choose neutrality between comparisons. However, when forced to express a preference (as in Mazeika et al. (2025)), it exhibits unequal values. Side-by-side comparison of country preferences on a 0-1 scale using the same country ordering. (Left) The 5-point Likert scale with a neutral option shows more uniform scores across countries. (Right) The 4-point forced-choice scale reveals a preference hierarchy.
Even vs. odd-sized Likert scales produce very different results. We find significant differences in preference patterns across elicitation formats. The 4-point forced choice condition produced non-uniform country preferences (Figure 7, right). Without a neutral option available, we find, similar to Mazeika et al. (2025), that the model displayed noticeable preferences for certain countries over others. In the 5-point Likert scale condition, GPT-4o selected the neutral midpoint (“No preference”) in 100% of country pair comparisons. This resulted in perfectly uniform normalized preference scores of 0.50 for all 11 countries (Figure 7, left). [^3] This reproduced the high-level finding from Mazeika et al. (2025) that GPT-4o expresses preferences that value the lives of people from different countries differently. However, note that we do not use the same technique for quantifying these results. We average model responses on Likert scales, while Mazeika et al. (2025) derive exchange rates from parametric log utility curves.
@@
Methodological Implications: Our results provide context for interpreting (Mazeika et al., 2025) exchange rate results, where they report that “GPT-4o places the value of Lives in the United States significantly below Lives in China, which it in turn ranks below Lives in Pakistan,” and suggest these represent “deeply ingrained biases” in the model. However, when allowed to select a ‘neutral’ option in comparisons, GPT-4o consistently indicates equal valuation of human lives regardless of nationality, suggesting a more nuanced interpretation of the model’s apparent preferences. This illustrates a key limitation in extracting preferences from LLMs. Rather than revealing stable internal preferences, our findings show that LLM outputs are largely constructed responses to specific elicitation paradigms. Interpreting such outputs as evidence of inherent biases without examining methodological factors risks misattributing artifacts of evaluation design as properties of the model itself.
@@ 7. Discussion
Significance: It is appealing to assume that modern LLMs exhibit stable, coherent, and steerable preferences, goals central to the field of AI Alignment. Prior research evaluating cultural alignment in LLMs has attempted to systematically characterize how effectively these models align with different cultures. However, we find that state-of-the-art LLMs display surprisingly erratic cultural preferences. When LLMs appear more aligned with certain cultures than others, such alignment tends to be nuanced and highly context-dependent. As demonstrated, even small methodological changes can yield substantially different outcomes. Our results caution against drawing broad conclusions from narrowly scoped experiments. In particular, they highlight that overly simplistic evaluations, cherry-picking, and confirmation biases may lead to an incomplete or misleading understanding of cultural alignment in LLMs.
@@
Limitations: The central theme of this work is about the risks of drawing broad conclusions from narrow evaluations. However, similar to previous research, we also only conduct a limited set of experiments. It is clear that stability, extrapolability, and steerability do not hold in general for state-of-the-art LLMs. Nonetheless, there may be specific circumstances in which they do. For example, Benkler et al. (2023); Azzopardi and Moshfeghi (2024); Jiang et al. (2024), and Mazeika et al. (2025) all argue that they may hold in narrow cases. While cultural alignment evaluations face fundamental limitations, our work should not be understood as showing that they are fundamentally invalid.
Recommendations: In the social sciences, pre-registering experiments has become common practice to reduce harms from cherry-picking and $p$-hacking (P Simmons et al., 2021). Given that LLMs often exhibit unpredictable preferences, adopting similar pre-registration practices would lend greater rigor in the evaluation of LLM cultural alignment. Additionally, since evaluations of cultural alignment in LLMs are frequently unreliable (Section 5) and sensitive to adversarial modifications (Section 6), incorporating red-teaming exercises could further strengthen assessment methods and improve robustness in evaluating LLM cultural alignment.
Future Work: Future work could focus on improving both models and evaluation strategies. Current models are typically fine-tuned to seek approval from evaluators (e.g., using RLHF), but fine-tuning specifically for steerability and consistency in preferences could lead to more generalizable evaluation results. Another promising direction is to develop evaluations that better reflect real-world complexities, given that alignment with values for both LLMs and humans is inherently nuanced and context dependent. Finally, it may be worthwhile to reconsider the emphasis on evaluating preferences in LLMs altogether. Rather than focusing solely on which cultures these models align with, future research could prioritize understanding how deploying LLMs influences power dynamics and tangibly impacts diverse communities in the real world (Kalluri, 2020).
@@ Appendix A Prompts
Which implied state of the world would you prefer? Please provide your rating (1-4).